# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 10 Affirmative Action in India
A common technical challenge encountered by U.S. Army officials integrating BRADSO incentives into the branching system and by Boston Public Schools (BPS) officials incorporating walk-zone priority into the student assignment plan was designing a mechanism that allocates some positions based on one criterion and others based on another. In both cases, officials found it relatively straightforward to define and implement policies using a single criterion, but introducing a second criterion for certain positions proved to be a formidable task.
This challenge extends beyond the Army and BPS. Mechanisms like the Army’s branching system and BPS’s school choice plan are technically complex, making it unrealistic to expect fully satisfactory designs from officials lacking expertise in formal fields, such as design economics or optimization. This underscores the importance of academic market designers expanding their focus beyond traditional economic principles, such as preference utilitarianism.
@@
The discovery of the erroneous implementation of the walk-zone priority at BPS from 2005 to 2013 (see 9.2), was the first indication that the subtle concept of order of precedence—the processing sequence of different types of positions—may be the source of significant unintended consequences in policy. Following this discovery, I systematically explored whether similar experiences existed worldwide where policymakers, lacking clarity on policy distinctions, intended one type of reserve policy but inadvertently implemented another.
@@
In this context, much of our focus in the next application is on a case where a flawed affirmative action procedure, mandated by the Supreme Court judgment in Anil Kumar Gupta vs. State of U.P. (1995), led to numerous lawsuits and disrupted recruitment processes in India for decades. Twenty-five years later, the Supreme Court addressed this design flaw in the subsequent judgment Saurav Yadav vs. State of U.P. (2020), but not before the issues and their remedies were thoroughly analyzed in Sönmez and Yenmez (2019, 2022).
This is another instance where life imitates science. Similar to the school choice reforms in England and Chicago discussed in Section 6.12, the amendment introduced in the Saurav Yadav vs. State of U.P. (2020) judgment provides external validation for minimalist market design. This landmark ruling aligns not only with Sönmez and Yenmez (2019, 2022) regarding the identified issues and their root causes but also with the proposed remedies outlined in these papers.
@@ 10.1 India’s Reservation System and the Quest for Social Justice
Affirmative action for various disadvantaged groups is embedded in the 1950 Constitution of India through articles 14-18.107107107This segment of the Constitution is referred to as the Equality Code in India. In practice, these protective policies are implemented through a reservation system that sets aside a certain fraction of positions in public jobs and seats at public schools for various protected groups. Naturally, the Constitution itself does not provide any rigorous formula to implement these policies. Nevertheless, in a landmark Supreme Court judgment Indra Sawhney vs. Union of India (1992), the justices formulated these provisions in a rigorous way, at least for the more basic versions of the reservation system.108108108Considered the main reference on the reservation system in India, this case is popularly known as the Mandal Commission case.
In the first part of this section, up to Section 10.7, following the analysis in Sönmez and Yenmez (2022), we focus on a version of the problem addressed in the Supreme Court judgments Anil Kumar Gupta vs. State of U.P. (1995) and Saurav Yadav vs. State of U.P. (2020). In this setting, a public institution allocates identical positions based on applicants’ merit scores and memberships in protected groups—a scenario common in practical applications across India. In Section 10.7, we extend the analysis to heterogeneous positions, as explored in Sönmez and Yenmez (2024), which represents another frequently encountered variant of the problem in the country.
@@
10.2 Vertical Reservations, Migration and the Over-and-Above Choice Rule
@@
From a technical point of view, the social identity-based vertical reservation system meant that no individual could be a member of multiple VR-protected groups. While VR protections are also provided to individuals in economic disability since 2019 with the enactment of the 103rd Amendment,109109109This more recent vertical category is called Economically Weak Segments (EWS). members of the earlier and social identity-based VR-protected groups are excluded from its scope,110110110This exclusion is highly controversial in India. See Sönmez and Ünver (2022) for a case study in minimalist market design that analyzes this exclusion in the 103rd Amendment. thereby still maintaining the non-overlapping structure of VR-protected groups in the country.
The defining property of the VR policy is formulated in paragraph 94 of Indra Sawhney vs. Union of India (1992):
@@
A subtle but important aspect of the above mandate is its reliance on a basic technical tool commonly referred to as migration in Indian legal terminology.111111111The concept of migration is also known as mobility in Indian legal terminology.
@@
However, because Indian legal terminology does not distinguish between “categories of individuals” and “categories of positions,” despite the mismatch between these two concepts, the terms “open” and “general” are used interchangeably in legal discussions. The concept of migration likely arises from this ambiguity in legal terminology. For example, when a member of the SC receives an open position based on merit, they are considered to have “migrated” from the SC to the general category. Relating this notion to the above mandate from Indra Sawhney vs. Union of India (1992), individuals who have migrated are those who “will be treated as open competition candidates.”
Effectively, designers of various reservation systems in the country, despite complex legal requirements, attempt to simplify the system by dividing the set of applicants into separate partitions and using SSD to allocate positions within each partition. Critically, however, when a natural partition does not exist, they resort to the notion of migration to artificially create one. In my view, reliance on this rudimentary technical tool, along with a second tool discussed in Section 10.3—(horizontal) adjustments—significantly exacerbates the legal and practical challenges of implementing India’s reservation system.
@@
10.3 Horizontal Reservations, Adjustments and the Minimum Guarantee Choice Rule
@@
In Sections 10.3 through 10.5, we focus on the simpler case of identical positions and non-overlapping HR protections, where each individual belongs to at most one HR-protected group. This scenario provides the clearest policy implications, as it aligns with the version of the problem reviewed in Saurav Yadav vs. State of U.P. (2020).
@@
In contrast to the VR policy, which is defined precisely for stand-alone implementation in Indra Sawhney vs. Union of India (1992), the description of the HR policy in paragraph 95 of the same judgment is notably less clear:
@@
This passage leaves unclear what exactly the “necessary adjustments” refer to and how they are to be made. The ambiguity surrounding these adjustments and the implementation of the HR policy was later clarified by the Supreme Court in Anil Kumar Gupta vs. State of U.P. (1995).112112112The formulation of the HR policy in Anil Kumar Gupta vs. State of U.P. (1995) is further clarified in a subsequent Supreme Court judgment, Rajesh Kumar Daria vs. Rajasthan Public Service Commission (2007). In this judgment, the Supreme Court further elaborated two key characteristics of the HR policy that distinguish it from the VR policy:
@@
Possibly to avoid complications arising from overlaps between “property rights” gained through these two types of affirmative action policies, the justices also recommended implementing the HR policy within each vertical category, including the open category.113113113This mode of HR policy is referred to as compartmentalized horizontal reservations in Anil Kumar Gupta vs. State of U.P. (1995). For example, if there are horizontal reservations for women, they are provided with preferential access for a fraction of the positions in each category (i.e., open, SC, ST, OBC, etc.) separately.
@@
Within this framework, the first characteristic outlined in Anil Kumar Gupta vs. State of U.P. (1995) clarified the legal notion of “adjustments”: it simply represented a minimum guarantee within each category.
@@
With this understanding, rather than relying on the rudimentary tool of adjustments, one can employ a generalization of the minimum guarantee choice rule—discussed in Section 9.1 for the case of a single protected category—to implement the HR policy within any specific category.
@@
In this generalization, given a set of applicants eligible for the category, the highest-merit beneficiaries of HR-protected groups are awarded HR-protected positions up to capacity in the first step. In the second step, any remaining positions—including unused HR-protected positions—are allocated to other individuals with the highest merit scores.
10.4 SCI-AKG vs. 2SMG: The Supreme Court’s Flawed Choice Rule and Its Minimalist Correction
For experts well-versed in formal modeling and analysis of reserve systems, integrating the HR policy within the VR policy is straightforward. With only the VR policy, the reservation system could be implemented by repeated application of the SSD for each category, starting with the open category. With both affirmative action policies, amending this procedure by simply replacing the SSD with the minimum guarantee choice rule for each category would do the trick. First introduced in Sönmez and Yenmez (2019, 2022), let’s refer to this procedure as the two-step minimum guarantee (2SMG) choice rule.
@@
However, justices are rarely accustomed to modeling with analytical methods and are, understandably, unfamiliar with the minimum guarantee choice rule—introduced in theoretical matching literature nearly two decades after Anil Kumar Gupta vs. State of U.P. (1995). Instead, the justices formalized and employed the notion of “adjustments” to operationalize the secondary HR policy alongside the primary VR policy.
@@
Their procedure begins by disregarding the HR policy in the first two steps and focusing solely on the VR policy. In the first step, open positions are tentatively allocated using SSD, followed by the allocation of VR-protected positions to the remaining eligible individuals, again using SSD, in the second step.114114114Note that these two steps together form the over-and-above choice rule.
@@
In the third step, any “necessary adjustments” are applied to incorporate the HR policy, if needed. For example, if the women’s minimum guarantee is unmet within a specific category, such as SC, at the end of the first two steps, the lowest-merit SC men tentatively admitted are replaced by the highest-merit unmatched SC women. This adjustment ensures the HR policy is accommodated within the SC category.
@@
This three-step outline of the Supreme Court’s procedure does not fully define a precise choice rule, as the adjustment process in the final step involves a particularly tricky aspect. Suppose that the minimum guarantee constraints are short by one woman both for the open category and also for the VR-protected category SC. Then, if there is an unmatched woman from SC, it is not clear whether she should replace a man from the open category for the adjustment process or a man from the SC category.
@@
The justices addressed this technical challenge by also partitioning the “adjusters” into distinct (i.e., non-overlapping) groups, and in an especially unfortunate mandate in Anil Kumar Gupta vs. State of U.P. (1995), they restricted access for open-category horizontal adjustments to members of the general category only. While, with this mandate, the set of individuals entitled to horizontal adjustments at each vertical category conveniently becomes uniquely defined for the last step of the procedure, it also plagues the resulting choice rule with highly consequential shortcomings. Following Sönmez and Yenmez (2022), let’s refer to this procedure, mandated in India between the years 1995 and 2020, as the Supreme Court of India–Anil Kumar Gupta (SCI-AKG) choice rule.
Before discussing the shortcomings of the SCI-AKG choice rule, I am inclined to speculate on what might have led the justices down this unfortunate path. In my view, the Supreme Court’s flawed choice design may have been shaped by the misguided synonymous use of the terms “open” category and “general” category, as well as the way the concept of “migration” is applied under the VR policy in the Indian legal framework.
@@
Recall that, in the first two steps of the SCI-AKG choice rule, positions are provisionally allocated using the over-and-above choice rule without accounting for the HR policy. Now, consider an unmatched SC woman at this stage. Why is she deemed ineligible for adjustments within the “open” category?
@@
Here is a hypothesis: Since she has not been provisionally awarded an “open” position (see Section 10.2), she has not yet “migrated” to the “general” category, according to Indian legal terminology, and remains a member of the SC category. Perhaps, then, still not being a member of the “general” category at the beginning of the third step, the misleading synonymous use of “open” and “general” categories might have led the justices to deem her ineligible for an adjustment within the “open” category.
The flaws of the SCI-AKG choice rule can be clearly demonstrated by relating it to the minimum guarantee choice rule. To facilitate this comparison, the following concept, rooted in the legal terminology of India used to describe various normative principles and reservation mechanisms, is useful.
@@
Leveraging the notion of meritorious reserved candidates, the SCI-AKG choice rule can alternatively be defined in two steps: First, open positions are allocated using the minimum guarantee choice rule, with eligibility restricted to members of the general category and meritorious reserved candidates. Next, reserved positions for each VR-protected category are allocated to the remaining eligible applicants using the minimum guarantee choice rule.
Observe that the only difference between the SCI-AKG and 2SMG choice rules is the set of applicants competing for open positions in the first step. Under the 2SMG choice rule, all applicants are considered for open positions. In contrast, the SCI-AKG choice rule restricts eligibility to members of the general category and meritorious reserved candidates. As a result, with the exception of meritorious reserved candidates, VR-protected individuals are excluded from accessing HR-protected positions in the open category.
This connection between the two choice rules sheds light on the root cause of the failures of the SCI-AKG choice rule.
Except for the recent EWS category, a member of a VR-protected category is not required to declare membership in their category.115115115This contrasts with the documentation needed to prove ineligibility for social identity (i.e., caste or tribal) based VR-protected categories, which is a prerequisite for claiming membership in the EWS category. This typically involves documentation of forward caste membership in India. By claiming their VR protections, an individual—unless a meritorious reserved candidate—loses their open category HR protections under the SCI-AKG choice rule. Thus, this mechanism violates a form of the incentive compatibility axiom, first formulated by Aygün and Bó (2021) in the context of affirmative action at Brazilian colleges.
More evident in practical applications than the failure of incentive compatibility, the SCI-AKG choice rule also violates a version of the no justified envy principle tailored to the current context:

### v4

@@ 10 Affirmative Action in India
A common technical challenge faced by U.S. Army officials integrating BRADSO incentives into the branching system and by Boston Public Schools (BPS) officials incorporating walk-zone priority into the student assignment plan was designing a mechanism that allocates some positions according to one criterion and others according to another. In both cases, officials found it relatively straightforward to define and implement policies based on a single criterion, but introducing a second criterion for certain positions proved to be a formidable task.
This challenge extends beyond the Army and BPS. Mechanisms like the Army’s branching system and BPS’s school choice plan are technically complex, making it unrealistic to expect fully satisfactory designs from officials lacking expertise in formal fields, such as economic design or optimization. This underscores the importance of academic market designers expanding their focus beyond traditional economic principles, such as preference utilitarianism.
@@
The discovery of the erroneous implementation of the walk-zone priority at BPS from 2005 to 2013 (see Section 9.2), was the first indication that the subtle concept of order of precedence—the processing sequence of different types of positions—may be the source of significant unintended consequences in policy. Following this discovery, I systematically explored whether similar experiences existed worldwide where policymakers, lacking clarity on policy distinctions, intended one type of reserve policy but inadvertently implemented another.
@@
In this context, much of our focus in the next application is on a case where a flawed affirmative action procedure, mandated by the Supreme Court judgment in Anil Kumar Gupta vs. State of U.P. (1995), led to numerous lawsuits and disrupted recruitment processes in India for decades. Twenty-five years later, the Supreme Court addressed this design flaw in the subsequent judgment Saurav Yadav vs. State of U.P. (2020), but not before the issues and their remedies were thoroughly analyzed in Sönmez and Yenmez (2019, 2022). A brief guide to the key judgments referenced in this section is presented in Table 5.
Table 5: Key judgments on India’s reservation system and their role in Section 10.
Case (Court, Year) Scope Role Section
Indra Sawhney v. Union of India (SC, 1992) VR and HR policies (general framework) Introduces VR, HR policies, and the “migration” concept; leaves joint VR and HR implementation unspecified. 10.110.3
Anil Kumar Gupta v. State of U.P. (SC, 1995) VR and HR policies (joint implementation, identical units) Introduces SCI–AKG via the concept of “adjustments”; restricts Open HR adjustments, leading to NJE failures. 10.310.5
Rajesh Kumar Daria v. RPSC (SC, 2007) VR and HR policies (joint implementation, identical units) Clarifies working of SCI–AKG and the notion of “adjustments” under Anil Kumar Gupta. 10.3
Saurav Yadav v. State of U.P. (SC, 2020) VR and HR policies (joint implementation, identical units) Mandates NJE; removes exclusion from Open HR adjustments; endorses 2SMG. 10.510.7
Tamannaben A. Desai v. Shital A. Nishar (Gujarat HC, 2020) VR and HR policies (joint implementation, identical units) Introduces procedure equivalent to 2SMG ; mandates in Gujarat. 10.5
Union of India v. Ramesh Ram (SC, 2010) VR and HR policies (joint implementation, heterogeneous units) Awards vacated Open seats after MRC migration to the general category (govt. jobs), violating NJE. 10.7
Tripurari Sharan v. Ranjit Kumar Yadav (SC, 2018) VR and HR policies (joint implementation, 10.7 units) Awards vacated Open seats after MRC migration to the category of the vacating MRC (medical colleges), violating NJE; contradicts Ramesh Ram. 10.7
  • Abbreviations: VR: Vertical Reservations; HR: Horizontal Reservations; NJE: No Justified Envy; SCI–AKG: Supreme Court of India–Anil Kumar Gupta; 2SMG: Two-Step Minimum Guarantee; MRC: Meritorious Reserved Candidate; SC: Supreme Court; HC: High Court.
This is another instance where life imitates science. Similar to the school choice reforms in England and Chicago discussed in Section 6.12, the amendment introduced in the Saurav Yadav (2020) judgment provides external validity for minimalist market design. This landmark ruling aligns not only with Sönmez and Yenmez (2019, 2022) regarding the identified issues and their root causes but also with the proposed remedies outlined in these papers.
@@ 10.1 India’s Reservation System and the Quest for Social Justice
Affirmative action for various disadvantaged groups is embedded in the 1950 Constitution of India through articles 14–18.120120120This segment of the Constitution is referred to as the Equality Code in India. In practice, these protective policies are implemented through a reservation system that sets aside a certain fraction of positions in public jobs and seats at public schools for various protected groups. Naturally, the Constitution itself does not provide any rigorous formula to implement these policies. Nevertheless, in a landmark Supreme Court judgment Indra Sawhney (1992), the justices formulated these provisions in a rigorous way, at least for the more basic versions of the reservation system.121121121Considered the main reference on the reservation system in India, this case is popularly known as the Mandal Commission case.
In the first part of this section (up to Section 10.7), following Sönmez and Yenmez (2022), we focus on a version of the problem addressed in the Supreme Court judgments Anil Kumar Gupta (1995) and Saurav Yadav (2020). In this setting, a public institution allocates identical positions based on applicants’ merit scores and membership in protected groups—a scenario common in practice across India. In Section 10.7, we extend the analysis to heterogeneous positions, as explored in Sönmez and Yenmez (2024), which represents another frequently encountered variant of the problem in the country.
@@
10.2 Vertical Reservations
@@
From a technical point of view, the social identity-based vertical reservation system meant that no individual could be a member of multiple VR-protected groups. While VR protections are also provided to individuals in economic disability since 2019 with the enactment of the 103rd Amendment,122122122This more recent vertical category is called Economically Weak Segments (EWS). members of the earlier and social identity-based VR-protected groups are excluded from its scope,123123123This exclusion is highly controversial in India. See Sönmez and Ünver (2022) for a case study in minimalist market design that analyzes this exclusion in the 103rd Amendment. thereby still maintaining the non-overlapping structure of VR-protected groups in the country.
The defining property of the VR policy is formulated in paragraph 94 of Indra Sawhney (1992):
@@
A subtle but important aspect of the above mandate is its reliance on a basic technical tool commonly referred to as migration in Indian legal terminology.124124124The concept of migration is also known as mobility in Indian legal terminology.
@@
However, because Indian legal terminology does not distinguish between “categories of individuals” and “categories of positions,” despite the mismatch between these two concepts, the terms “open” and “general” are used interchangeably in legal discussions. The concept of migration likely arises from this ambiguity in legal terminology. For example, when a member of the SC receives an open position based on merit, they are considered to have “migrated” from the SC to the general category. Relating this notion to the above mandate from Indra Sawhney (1992), individuals who have migrated are those who “will be treated as open competition candidates.”
Effectively, designers of various reservation systems in the country, despite complex legal requirements, attempt to simplify the system by (1) dividing the set of applicants into separate partitions and (2) using SSD to allocate positions within each partition. Critically, however, when a natural partition does not exist, they resort to the notion of migration to artificially create one. In my view, reliance on this rudimentary technical tool, together with a second tool discussed in Section 10.3—(horizontal) adjustments—significantly exacerbates the legal and practical challenges of implementing India’s reservation system.
@@
10.3 Horizontal Reservations
@@
In Sections 10.3 through 10.5, we focus on the simpler case of identical positions and non-overlapping HR protections, where each individual belongs to at most one HR-protected group. This scenario provides the clearest policy implications, as it aligns with the version of the problem reviewed in Saurav Yadav (2020).
@@
In contrast to the VR policy, which is defined precisely for stand-alone implementation in Indra Sawhney (1992), the description of the HR policy in paragraph 95 of the same judgment is notably less clear:
@@
This passage leaves unclear what exactly the “necessary adjustments” refer to and how they are to be made. The ambiguity surrounding these adjustments and the implementation of the HR policy was later clarified by the Supreme Court in Anil Kumar Gupta (1995).125125125The formulation of the HR policy in Anil Kumar Gupta (1995) is further clarified in a subsequent Supreme Court judgment, Rajesh Kumar Daria vs. Rajasthan Public Service Commission (2007). In this judgment, the Supreme Court further elaborated two key characteristics of the HR policy that distinguish it from the VR policy:
@@
Possibly to avoid complications arising from overlaps between “property rights” gained through these two types of affirmative action policies, the justices also recommended implementing the HR policy within each vertical category, including the open category.126126126This mode of HR policy is referred to as compartmentalized horizontal reservations in Anil Kumar Gupta (1995). For example, if there are horizontal reservations for women, they are provided with preferential access for a fraction of the positions in each category (i.e., open, SC, ST, OBC, etc.) separately.
@@
Within this framework, the first characteristic outlined in Anil Kumar Gupta (1995) clarified the legal notion of “adjustments” in Indra Sawhney (1992) in defining horizontal reservations: it simply represented a minimum guarantee within each category.
@@
With this understanding, rather than relying on the rudimentary tool of adjustments, one can employ a computationally more efficient algorithm—the procedure used to introduce the minimum guarantee choice rule in Section 9.1—to implement the HR policy within any given category.
@@
In this algorithm, given a set of applicants eligible for the category, the highest-merit beneficiaries of HR-protected groups are awarded HR-protected positions up to capacity in the first step. In the second step, any remaining positions—including unused HR-protected positions—are allocated to other individuals with the highest merit scores.
10.4 SCI–AKG vs. 2SMG: The Supreme Court’s Flawed Choice Rule and Its Minimalist Correction
Building directly on the minimum guarantee choice rule, experts well-versed in formal modeling and analysis of reserve systems will find that integrating the HR policy within the VR policy is straightforward.
Two-step Minimum Guarantee (2SMG) Choice Rule.
With only the VR policy, the reservation system is implemented by repeated application of the SSD for each category, starting with the open category. With both VR and HR policies, this procedure can be amended by simply replacing the SSD with the minimum guarantee choice rule for each category. Moreover, when VR-protected categories do not overlap—as in India—all of them can be processed simultaneously after the open category, thereby justifying the name of the two-step minimum guarantee choice rule.
Let’s illustrate 2SMG with an example, using a continuum of positions and individuals to simplify its presentation. We first describe the setting, which we will also use later to explain the Supreme Court’s choice rule in Anil Kumar Gupta (1995).
Example 12.
There is one VR-protected group R and one HR-protected group W (women). There is a measure of 360 open positions, with a minimum guarantee of 150 for women. There is a measure of 60 VR-protected positions set aside for category R, with a minimum guarantee of 18 for category-R women. There is a measure of 300 general-category men, 300 category-R men, 100 general-category women, and 200 category-R women, each with scores uniformly distributed between 0 and 100. Panel (a) of Figure 26 represents this economy.
Figure 26: Execution of the 2SMG choice rule under its standard formulation. Panel (a) illustrates the continuum economy in Example 12. Building on Example 26, Panel (b) depicts how the 2SMG choice rule operates for this economy. Abbreviations: P: Positions; W: Women; R: Vertically Reserved; $M^{G}$: General-category men; $M^{R}$ Reserved-category men; $W^{G}$: General-category women; $W^{R}$: Reserved-category women.
We now turn to the implementation of the 2SMG choice rule in this setting, depicted in Panel (b) of Figure 26.
Example 12a. The process starts by allocating the 360 measure of open positions using the minimum guarantee choice rule. First, a measure of 150 is allocated to the highest-scoring women from both categories: a measure of 50 general-category women and a measure of 100 category-R women with scores 50 and higher. Next, the remaining 210 measure of open positions is allocated to the highest-scoring remaining individuals. Since there is more than 210 measure of men with scores higher than 50 (the upper bound of scores for the remaining women), all of these units are awarded to men: 105 measure to general-category men and 105 measure to category-R men with scores 65 and higher.
The process then continues by allocating the 60 measure of VR-protected positions set aside for category R, again using the minimum guarantee choice rule. First, a measure of 18 is allocated to the highest-scoring remaining category-R women, namely those with scores between 41 and 50. Next, the remaining 42 measure of category-R positions is allocated to the highest-scoring remaining individuals from category R. Since there is more than 42 measure of category-R men with scores higher than 41 (the upper bound of scores for the remaining category-R women), these units are awarded to category-R men with scores between 51 and 65.
@@
As natural as 2SMG is as a generalization of the over-and-above choice rule to integrate the secondary HR policy with the primary VR policy, justices—rarely accustomed to modeling with analytical methods and, understandably, unfamiliar with the minimum guarantee choice rule, introduced in the theoretical matching literature nearly two decades after Anil Kumar Gupta (1995). Instead, they relied on the notion of “adjustments” to operationalize the secondary HR policy alongside the primary VR policy.
SCI–AKG Choice Rule.
@@
The Supreme Court’s three-step procedure begins by ignoring the HR policy in the first two steps and focusing solely on the VR policy. In the first step, open positions are tentatively allocated using SSD. In the second step, the remaining VR-protected positions are allocated, again using SSD. Together, these two steps form our familiar over-and-above choice rule.
@@
In the third step, “necessary adjustments” are applied to incorporate the HR policy when needed. For example, if the women’s minimum guarantee is unmet within a specific category, such as SC, the lowest-merit SC men tentatively admitted are replaced by the highest-merit unmatched SC women. This adjustment helps ensure that the HR policy is accommodated within the SC category.127127127Since adjustments for the open category may require revising the allocation of VR-protected positions in Step 2, it is computationally more efficient to delay this step until adjustments for the open category have been finalized.
@@
Yet even this three-step description does not fully specify a precise choice rule, because the adjustment process in the final step involves a particularly tricky ambiguity. Suppose that the minimum guarantee constraints are short by one woman both for the open category and for the SC category. If there is an unmatched SC woman, it is unclear whether she should replace a man from the open category or one from SC itself.
@@
The justices attempted to resolve this difficulty by partitioning the “adjusters” into distinct (i.e., non-overlapping) groups. In an especially unfortunate mandate in Anil Kumar Gupta (1995), they further restricted access for open-category horizontal adjustments to members of the general category only. While this mandate conveniently ensured that the set of individuals entitled to horizontal adjustments at each vertical category became uniquely defined for the last step of the procedure, it also plagued the resulting choice rule with highly consequential shortcomings. Following Sönmez and Yenmez (2022), we refer to this procedure, mandated in India between 1995 and 2020, as the Supreme Court of India–Anil Kumar Gupta (SCI–AKG) choice rule.
We next illustrate the SCI–AKG choice rule using the economy in Example 12, shown in Panel (a) of Figure 26.128128128Strictly speaking, this illustration employs the slight variant mentioned in Footnote 127.
Example 12b. The process starts by tentatively allocating the 360 measure of open positions to the highest-scoring individuals across all groups: a measure of 120 general-category men, a measure of 120 category-R men, a measure of 40 general-category women, and a measure of 80 category-R women with scores of 60 and higher (see Panel (a) of Figure 27). Since women received only a measure of 120 at this stage—below the minimum guarantee of 150—a measure of 30 of the lowest-scoring men tentatively admitted must be dropped as part of the horizontal adjustment process for the open category. This results in the removal of a measure of 15 general-category men and a measure of 15 category-R men, those with scores between 60 and 65 (see Panel (b) of Figure 27, where this “adjustment” is indicated with A).
To complete the adjustment process for the open category, this measure of 30 positions is awarded to general-category women, since category-R women are deemed ineligible for the open-category adjustment. Consequently, a measure of 30 general-category women with scores between 30 and 60 receive these adjusted units (see Panel (c) of Figure 27, where this “adjustment” is indicated with A).
The process then continues by allocating the 60 measure of VR-protected positions set aside for category R to the highest-scoring members of this category. Since no remaining category-R women have scores above 60, the first measure of 15 goes to category-R men with scores between 60 and 65, before the process begins awarding positions to category-R women. At this point, the remaining measure of 45 is split between a measure of 27 category-R men and a measure of 18 category-R women with scores between 51 and 60. Since the minimum guarantee of a measure of 18 VR-protected units for women is met, no further adjustment is necessary, finalizing the process (see Panel (d) of Figure 27).
Importantly, since all general-category women with scores of 30 and higher received positions, category-R women with scores between 30 and 51—who remain unassigned—have justified envy toward some of these women.
Figure 27: Execution of the SCI–AKG choice rule. For the continuum economy in Example 12, given in Panel (a) of Figure 26, Panel (a) here illustrates the allocation of open positions under the VR policy only. Panels (b) and (c) show the subsequent adjustment process (indicated with A) for the open category to accommodate the minimum guarantee for women. In Panel (b), a measure of 30 lowest-scoring men (15 from each category) lose their tentative assignments from Panel (a), which in Panel (c) are awarded exclusively to the highest-scoring unmatched women from the general category. Panel (d) depicts the allocation of VR-protected positions. Abbreviations: P: Positions; W: Women; R: Vertically Reserved; $M^{G}$: General-category men; $M^{R}$ Reserved-category men; $W^{G}$: General-category women; $W^{R}$: Reserved-category women.
Before further discussing the shortcomings of the SCI–AKG choice rule, as illustrated at the end of Example 10.4, I am inclined to speculate on what might have led the justices to this unfortunate design, as this will help us identify the root cause of the problem.
@@
In my view, the Supreme Court’s flawed choice rule may have been shaped by the misguided conflation of the terms “open” category and “general” category, as well as by the way the concept of “migration” is applied under the VR policy in the Indian legal framework. Recall that, in the first two steps of the SCI–AKG choice rule, positions are provisionally allocated using the over-and-above choice rule without accounting for the HR policy. Now, consider an unmatched VR-protected woman at this stage. Why is she deemed ineligible for women’s horizontal adjustments within the “open” category?
@@
One possible explanation is as follows: Since she has not been provisionally awarded an “open” position (see Section 10.2), she has not yet “migrated” to the “general” category, according to Indian legal terminology, and remains a member of her VR-protected category. Perhaps, then, because she was not yet considered a member of the “general” category at the beginning of the third step, the misleading conflation of “open” and “general” categories might have led the justices to deem her ineligible for an adjustment within the open category.
To understand why SCI–AKG fails, it is instructive to recall a notion from Indian legal terminology that relates this design to 2SMG.
@@
Alternative Formulation for SCI–AKG.
The justices’ design can be reformulated in terms of meritorious reserved candidates, using repeated applications of the minimum guarantee choice rule. In the first step, open positions are allocated using the minimum guarantee choice rule, with eligibility restricted to members of the general category and meritorious reserved candidates. In the second step, reserved positions for each VR-protected category are allocated to the remaining eligible applicants, again using the minimum guarantee choice rule.
The only difference between this formulation of SCI–AKG and the 2SMG choice rule lies in the set of applicants competing for open positions in the first step. Under the 2SMG choice rule, all applicants are considered. By contrast, with the exception of the meritorious reserved candidates, the SCI–AKG choice rule restricts eligibility to members of the general category. As a result, VR-protected individuals other than meritorious reserved candidates are excluded from accessing HR-protected positions in the open category.
This connection exposes not only the shortcomings of the SCI–AKG choice rule but also their root cause.
Except for the recent EWS category, a member of a VR-protected category is not required to declare membership in their category.129129129This contrasts with the documentation needed to prove ineligibility for social identity (i.e., caste or tribal) based VR-protected categories, which is a prerequisite for claiming membership in the EWS category. This typically involves documentation of forward caste membership in India. By claiming their VR protections, an individual—unless they are a meritorious reserved candidate—loses their open category HR protections under the SCI–AKG choice rule. Thus, this mechanism violates a form of the incentive compatibility axiom, first formulated by Aygün and Bó (2021) in the context of affirmative action at Brazilian colleges.
More evident in practical applications than the failure of incentive compatibility, as we illustrated in Example 10.4, the SCI–AKG choice rule also violates a version of the no justified envy principle tailored to the current context: